Suppose Rocky Brands has earnings per share of ​$2.46 and EBITDA of ​$29.9 million. The firm also has 5.25 million shares outstanding and debt of ​$135 million​ (net of​ cash). You believe Deckers Outdoor Corporation is comparable to Rocky Brands in terms of its underlying​ business, but Deckers has no debt. If Deckers has a​ P/E of 13.2 and an enterprise value to EBITDA multiple of 7.2, estimate the value of Rocky Brands stock using both multiples. Which estimate is likely to be more​ accurate?
Answer #1

Solution:-

Rocky brand's valuation using EV/EBITDA ratio:

Rocky Brand's EBITDA= $29.9m

Comparable EV/EBITDA ratio of Deckers is 7.2 which if used for Rocky brands, its EV is as follows:

Rocky Brand's Enterprise value= EBITDA*7.2= $29.9m*7.2= $215.28

Now, we know that the Enterprise value has the following formula:

Enterprise value= Market value of equity + net debt

215.28= Market cap of equity + 135

Market cap of equity= 215.28-135= $80.28m

Value per share= $80.28m/No. of shares= $80.28m/5.25= $15.29 per share

Rocky brand's valuation using PE ratio:

Rocky Brand's EPS= $2.46 per share

Comparable PE ratio of Deckers is 13.2 times which if used for Rocky brand would yield the following market value per share

EPS*PE ratio= Value per share

$2.46*13.2= Value per share

Value per share= $32.47 per share

Which estimate is likely to be more accurate:

Though PE ratio is the more direct measure of a firm's equity valuation, however EV/EBITDA multiple is better suited to compare peers that have different capital structures. This is due to the fact that EV/EBITDA is based on operating income for both debt and equity and therefore negates the impact of varying capital structures among comparable peers. Due to the above reason, the value per share of $15.29 as arrived using EV/EBITDA multiple is likely to be more accurate.